# Multi-Robot Navigation in Social Mini-Games: Definitions, Taxonomy, and Algorithms

### v3

@@ Abstract
Multi-robot navigation (MRN) is an active and essential area of research in robotics and artificial intelligence with wide-ranging applications such as autonomous driving, warehouse logistics, swarms, delivery, service, and personalized home robots [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. MRN builds upon the challenges of single-robot navigation by introducing the complexities of inter-robot, human-robot interactions, and collective decision-making. In multi-robot systems, the algorithms must account for the intentions, movements, and potential conflicts among multiple robots, including humans, making the navigation problem significantly more complex [11, 12, 13].
@@
In recent years, the research field of MRN has started to coincide with the momentum of robot deployments in shared human spaces, largely driven by advances in data, compute, and foundation models. As a consequence, research has flourished in MRN, and more specifically, MRN in constrained environments that arise in shared human spaces. In Figure 1, we highlight two types of MRN in constrained scenarios. On the left, we depict Social Mini-Games (SMGs) [14, 16] on the right we depict “general MRN”. A key distinguishing factor between these scenarios is the notion of agency–that is, the capacity of an agent to take independent actions and influence the environment around them. SMGs exhibit high agency where small actions can significantly change the interaction dynamics, coordination requirements, and potential for deadlocks. General MRN, on the other hand, exhibit low to medium agency. High-agency SMGs have several unique challenges. First, without explicit scheduling, robots result in deadlocks, collisions, or non-smooth trajectories [17, 18]. Second, humans can avoid collisions and deadlocks without having to deviate too much from their nominal speed or trajectory. For instance, when two individuals go through a doorway together, one person modulates their velocity by just enough to enable the other person to pass through first, while still adhering closely to their preferred speed. This type of behavior presents a significant challenge for robots, especially in high-agency SMGs. Classical navigation approaches typically fail to cover all the aspects needed for SMGs [14]. So, new techniques, ideas, and systems are being investigated, often with roots deep within the larger parent field of MRN. This fact presents the crux of the issue: researchers studying SMG solvers from a MAPF perspective will build off of MAPF research, MARL-based SMG solvers will form their roots from the MARL community, and so on. Other communities build upon game theory and optimization. Each field has its own set of assumptions, limitations, objectives, and techniques. As a result, MRN in SMG research has fractured into these sub-communities without a clear, unified formulation. This categorization into sub-communities presents an entry barrier to new researchers in this field by making it difficult to navigate through and understand existing literature and to establish appropriate baselines for comparison. In addition, this makes it difficult for practitioners to find papers and solutions relevant to their concrete application. This article aims to address this growing challenge by introducing a unified taxonomy to describe MRN in SMG problems, and by establishing common assumptions, properties, and measures for evaluating MRN in SMG algorithms. The unified taxonomy we present in this paper is our attempt to classify the currently studied SMG solver variants. We hope this terminology will serve as a common ground for future researchers, and will be used by them to describe their contributions succinctly and accurately.
@@ 1.1 Differentiation from Existing Surveys
Our survey focuses on multi-robot navigation in social mini-games, and thus belongs to the broader field of social robot navigation. Social robot navigation, however, is vast and we will not attempt to summarize it; instead, we refer to many recent surveys on social navigation [19, 20, 21, 22, 23, 24, 25, 26, 27, 28]. Among these, [24] focuses on evaluating social robot navigation algorithms, reviewing 177 recent papers to gather evaluation methods, scenarios, datasets, and metrics, using their findings to discuss shortcomings of existing research and to make recommendations on how to properly evaluate social navigation research. Another recent survey by Mavrogiannis et al. [25] focuses on the core challenges of social navigation with respect to different algorithms, human behavior models, and evaluation, but they focus on the general navigation case. Wang et al. [28] proposes new metrics evaluating the principles defined in [19], such as comfort, naturalness, and sociability. And finally, datasets [29, 30] and benchmarks [31, 32] have enabled researchers to compare algorithms, revealing limitations of existing solutions, and illuminating promising new directions. However, all of the surveys discussed above on various aspects of social robot navigation focus on robots in sparse or dense crowds, but not social mini-games. We showcase the difference in Figure 1. Perhaps the closest survey that mirrors the same objective as with this survey (that is, attempting to unify a growing, but fractured, research area) is the seminal survey on MAPF (“Multi-Agent Pathfinding: Definitions, Variants, and Benchmarks”) by Stern et al. [33]. Here, the authors present a definition for MAPF, describe the different variants of it, and introduce two benchmarks for proper evaluation. This paper was influential in anchoring and streamlining the future research in MAPF. We envision a similar success via our survey. The overall structure of this survey is given in Fig. 2. We first define the SMG problem in Sec. 2, highlight its various scenarios where it appears, and how to evaluate navigation approaches in SMGs. Then in Sec. 3, we introduce the taxonomy, properties, and algorithmic categories for research on SMG solvers. Then in Sec. 4, we benchmark and compare some recent state of the art SMG solvers on the scenarios introduced earlier. Finally in Sec. 5, we conclude by discussing general trends, open problems, and future directions.
@@ 2 Social Mini-Games
2.1 SMGs versus General MRN
SMGs are a special class of MRN scenarios. Every SMG is an instance of an MRN, but the other way around is not true. Specifically, SMGs are high-agency MRN scenarios. Here, agents are not only moving in close proximity to other agents, but the key distinguishing characteristic of SMGs is that, unless some agent changes its course or takes a different action than its optimally decided one, there will be an undesirable outcome. That is what is meant by high agency–each agent is significantly impacting other agents such that they need to coordinate with each other to survive. Examples of SMGs are shown in Figure 3 and discussed in detail in the following Section. But from a glance, it can be observed that each of these scenarios is an SMG. Consider for example the doorway scenario: unless either of the green or orange agent alters its course, they will collide.
We next give a mathematical definition of SMGs. It says that if we assume each agent has one or more individual optimal trajectory (for example, in the doorway scenarios, the optimal trajectory for both the green and the orange agent would be a straight line path towards the red cross.), then if all the individual optimal trajectories intersect for all agents within some time interval, then that scenario is an SMG.
Beyond saying that SMGs are a subset of general MRN, it is important to highlight the exact characteristics that distinguish the two types of scenarios. We give such a list below.
@@ 2.2 Common SMG Scenarios
In this Section, we briefly list a non-exhaustive set of SMG scenarios that occur in everyday life. These scenarios are compiled from a combination of prior social navigation literature [34, 35] and robotics literature [36, 37]. Each canonical scenario activates a distinct subset of SMG characteristics. While not exhaustive, they are chosen to cover the most frequently encountered topologies in real life such as doorways, intersections, and blind corners that expose different deadlock mechanisms, visibility challenges, and coordination demands. Additional scenarios, as and when they are discovered and analyzed, will be added to our accompanying software library.
@@ 2.3 Evaluation Metrics
The typical metrics for evaluating general MRN are the following: Avg , Makespan Ratio, and Path Deviation. We chose these metrics for benchmarking MRN for SMGs because they provide a comprehensive view of each method’s performance. The Avg metric helps measure movement smoothness, the Makespan Ratio compares efficiency between agents, Path Deviation assesses navigation accuracy, and the Deadlock Rate evaluates how often the agents are stuck in deadlocks. Together, these metrics ensure that the path planners can handle real-world challenges effectively, navigating efficiently, smoothly, and reliably while avoiding deadlocks. This thorough evaluation is essential for developing robust multi-agent systems in complex environments. Although existing metrics for general MRN remain essential, in an SMG, agents have high agency. We therefore also report the following metrics: flow rate, fairness (system-level performance), and invasiveness (externalities imposed on others). The following sections discuss each of them in more detail:
@@
Avg (Average Change in Velocity): This metric measures how much an agent’s speed changes on average at each time step during its movement. It shows how often and how much the agent adjusts its speed in response to its surroundings and other agents. It is given as: , where is summed over the entire episode over which the metric is evaluated. A lower Avg means smoother and more consistent movement, with fewer disruptions or sudden changes. A higher Avg means the agent is often changing its speed, which could indicate navigation challenges or the need for constant adjustment in a changing environment.
Makespan Ratio: The Makespan Ratio is the comparison of the time taken by one (concerned) agent to reach its goal (TTG) to the time taken by another (fastest) agent to reach the same goal, i.e. , where is the time-to-goal for the agent. It gives a relative measure of how efficiently the agents navigate. A Makespan Ratio close to 1 means that the agents take about the same amount of time to reach their goals, indicating balanced and fair navigation strategies. If the ratio is far from 1, it might show differences in navigation times, possibly due to bottlenecks, poor path choices, or other navigation challenges.
@@
Path Deviation: Path Deviation measures how much an agent’s path differs from a reference or ideal path. It uses the Hausdorff distance, which calculates the greatest distance from a point in one path to the closest point in the other path, i.e.
(1) where and are the sets of points in the actual path taken by the agent, and the set of points in the nominal path, respectively. Note that the metric in Eq 1 is purposefully kept broad so that it can be adapted to other deviation metrics by the user.
@@
In this case, the deviation is measured against the ideal scenario where only one agent moves towards the goal without any interactions or obstacles, defined as the nominal agent scenario. A low Path Deviation means the agent’s actual path is very close to the optimal path, indicating efficient navigation. A high Path Deviation means the agent took significant detours or deviations, possibly due to interactions with other agents, obstacles, avoiding or resolving deadlocks in the way or irregular control generation leading to poor navigation.
Flow Rate: Flow-rate measures the flow through a gap or a door-way where multiple agents are trying to get through. It is defined as: , where is the number of agents, is the make-span (in sec.), while is the gap width (in unit meters) [14]. Flow-rate gives a metric quantifying the overall ‘volume of agents‘ successfully navigating a constrained space.
Fairness: Let is a risk function that determines preference over trajectories, , where represents the private valuation of an agent , quantifying its preference or risk associated with a given trajectory. To formalize fairness at the system level, we define a global cost , also known as social welfare and represents the individual agent’s local cost. For mathematical simplicity, we will now express this in terms of rewards instead of costs: .
Thus, a fair outcome is defined as a set of control inputs or trajectories such that (1) Each agent is content with their strategy/policy and outcome, meaning their individual reward is maximized given their chosen strategy: and (2) the total social welfare, accounting for all agents’ preferences, is maximized at the system level, that is, .
Invasiveness Score (IS): Measures the control deviation an agent induces on others:
where is agent ’s control under the full multi-agent policy and is the counterfactual control without agent . A higher indicates more invasive behavior.

### v4

@@ Abstract
Multi-robot navigation (MRN) is an active and essential area of research in robotics and artificial intelligence with wide-ranging applications such as autonomous driving, warehouse logistics, swarms, delivery, service, and personalized home robots [drew2021multi, rasheed2022review, loizou2014multi, nathan2011review, orr2023multi, antonyshyn2023multiple, chen2021decentralized, raibail2022decentralized, huang2019collision, patwardhan2022distributing]. MRN builds upon the challenges of single-robot navigation by introducing the complexities of inter-robot, human-robot interactions, and collective decision-making. In multi-robot systems, the algorithms must account for the intentions, movements, and potential conflicts among multiple robots, including humans, making the navigation problem significantly more complex [mavrogiannis2017socially, fan2020distributed, alzetta2020time].
@@
In recent years, the research field of MRN has started to coincide with the momentum of robot deployments in shared human spaces, largely driven by advances in data, compute, and foundation models. As a consequence, research has flourished in MRN, and more specifically, MRN in constrained environments that arise in shared human spaces. In Figure 1, we highlight two types of MRN in constrained scenarios. On the left, we depict Social Mini-Games (SMGs) [chandra2025deadlock, chandra2024towards] on the right we depict “general MRN”. A key distinguishing factor between these scenarios is the notion of agency–that is, the capacity of an agent to take independent actions and influence the environment around them. SMGs exhibit high agency where small actions can significantly change the interaction dynamics, coordination requirements, and potential for deadlocks. General MRN, on the other hand, exhibit low to medium agency. High-agency SMGs have several unique challenges. First, without explicit scheduling, robots result in deadlocks, collisions, or non-smooth trajectories [kandhasamy2020scalable, qian2015decentralized]. Second, humans can avoid collisions and deadlocks without having to deviate too much from their nominal speed or trajectory. For instance, when two individuals go through a doorway together, one person modulates their velocity by just enough to enable the other person to pass through first, while still adhering closely to their preferred speed. This type of behavior presents a significant challenge for robots, especially in high-agency SMGs. Classical navigation approaches typically fail to cover all the aspects needed for SMGs [chandra2025deadlock]. So, new techniques, ideas, and systems are being investigated, often with roots deep within the larger parent field of MRN. This fact presents the crux of the issue: researchers studying SMG solvers from a MAPF perspective will build off of MAPF research, MARL-based SMG solvers will form their roots from the MARL community, and so on. Other communities build upon game theory and optimization. Each field has its own set of assumptions, limitations, objectives, and techniques. As a result, MRN in SMG research has fractured into these sub-communities without a clear, unified formulation. This categorization into sub-communities presents an entry barrier to new researchers in this field by making it difficult to navigate through and understand existing literature and to establish appropriate baselines for comparison. In addition, this makes it difficult for practitioners to find papers and solutions relevant to their concrete application. This article aims to address this growing challenge by introducing a unified taxonomy to describe MRN in SMG problems, and by establishing common assumptions, properties, and measures for evaluating MRN in SMG algorithms. The unified taxonomy we present in this paper is our attempt to classify the currently studied SMG solver variants. We hope this terminology will serve as a common ground for future researchers, and will be used by them to describe their contributions succinctly and accurately.
@@ 1.1 Differentiation from Existing Surveys
Our survey focuses on multi-robot navigation in social mini-games, and thus belongs to the broader field of social robot navigation. Social robot navigation, however, is vast and we will not attempt to summarize it; instead, we refer to many recent surveys on social navigation [kruse2013human, rios2015proxemics, chik2016review, charalampous2017recent, cheng2018autonomous, gao2021evaluation, mavrogiannis2021core, mirsky2021prevention, moller2021survey, wang2022metrics]. Among these, [gao2021evaluation] focuses on evaluating social robot navigation algorithms, reviewing 177 recent papers to gather evaluation methods, scenarios, datasets, and metrics, using their findings to discuss shortcomings of existing research and to make recommendations on how to properly evaluate social navigation research. Another recent survey by Mavrogiannis et al. [mavrogiannis2021core] focuses on the core challenges of social navigation with respect to different algorithms, human behavior models, and evaluation, but they focus on the general navigation case. Wang et al. [wang2022metrics] proposes new metrics evaluating the principles defined in [kruse2013human], such as comfort, naturalness, and sociability. And finally, datasets [scanddataset, martin2019jrdb] and benchmarks [kastner2022arena, sprague2023socialgym] have enabled researchers to compare algorithms, revealing limitations of existing solutions, and illuminating promising new directions. However, all of the surveys discussed above on various aspects of social robot navigation focus on robots in sparse or dense crowds, but not social mini-games. We showcase the difference in Figure 1.
Perhaps the closest survey that mirrors the same objective as with this survey (that is, attempting to unify a growing, but fractured, research area) is the seminal survey on MAPF (“Multi-Agent Pathfinding: Definitions, Variants, and Benchmarks”) by Stern et al. [stern2019multi]. Here, the authors present a definition for MAPF, describe the different variants of it, and introduce two benchmarks for proper evaluation. This paper was influential in anchoring and streamlining the future research in MAPF. We envision a similar success via our survey. The overall structure of this survey is given in Fig. 2. We first define the SMG problem in Sec. 2, highlight its various scenarios where it appears, and how to evaluate navigation approaches in SMGs. Then in Sec. 3, we introduce the taxonomy, properties, and algorithmic categories for research on SMG solvers. Then in Sec. 4, we benchmark and compare some recent state of the art SMG solvers on the scenarios introduced earlier. Finally in Sec. 5, we conclude by discussing general trends, open problems, and future directions.
@@ 2 Social Mini-Games
2.1 MRN Preliminaries
We first begin by defining a general MRN as a partially observable stochastic game: where denotes the number of robots. Hereafter, will refer to the index of a robot and appear as a superscript whereas will refer to the current time-step and appear as a subscript. The general state space (e.g. SE(2), SE(3), etc.) is continuous; the robot at time has a state . A state consists of both visible parameters (e.g. current position, linear and angular velocity and hidden (to other agents) parameters which could refer to the internal state of the robot such as preferred speed, preferred heading, etc. We denote the set of observable parameters as . On arriving at a current state , each robot generates a local observation, , via , where , the set of robots detected by ’s sensors. Over a finite horizon , each robot is initialized with a start state , a goal state where and denote subsets of containing the initial and final states. The transition function is given by , where is the continuous control space for robot representing the set of admissible inputs for . A discrete trajectory is specified as and its corresponding input sequence is denoted by . Robots follow the discrete-time control-affine system, . Let be a policy that maps the observation or state history of agent to an action
We denote by as the set of preferred trajectories for robot that solve the two-point boundary value problem. A preferred trajectory, as defined by existing methods [impc, orca, nh-orca], refers to a collision-free path a robot would follow in the absence of dynamic or static obstacles and is generated via a default planner according to some predefined criteria such as shortest path, minimum time, etc. A collision is defined as follows. Let represent the space occupied by robot (as a subset of the state space ) at any time which can be approximated by the convex hull of a finite number of points that determine the boundary of the robot (e.g., vertices of a polytopic set). Then, robots and are said to collide at time if . Each robot has a running cost that assigns a cost to a state-input pair at each time step based on user-defined variables e.g. distance of the robots current position from the goal, change in the control across subsequent time steps, and distance between the robots preferred and actual paths.
2.2 From General MRN to SMGs
SMGs are a special class of MRN scenarios. Every SMG is an instance of an MRN, but the other way around is not true. Informally, SMGs arise when all the optimal or preferred trajectories of two or more agents collide within a time interval. We make this formal via the definition below.
In practice, we approximate collision by overlap of convex hulls inflated by a safety margin. We depict several examples and non-examples of social mini-games in Figure 3. The first scenario can be characterized as a social mini-game due to the conflicting preferred trajectories of agents and within a specific time interval , where the duration . In this case, the agents’ trajectories intersect, generating a conflict. However, the second and third scenarios, in contrast, do not qualify as social mini-games since no conflicts arise between the agents. In the second scenario, there is no common time duration where agents intersect each other, and their trajectories remain independent in time. In the third scenario, agent possesses an alternative preferred trajectory that avoids conflicts during any time duration, allowing for a seamless transition to a conflict-free path.
SMGs are high-agency MRN scenarios. Here, agents are not only moving in close proximity to other agents, but the key distinguishing characteristic of SMGs is that, unless some agent changes its course or takes a different action than its optimally decided one, there will be an undesirable outcome. That is what is meant by high agency–each agent is significantly impacting other agents such that they need to coordinate with each other to survive. Examples of SMGs are shown in Figure 5 and discussed in detail in the following Section. But from a glance, it can be observed that each of these scenarios is an SMG. Consider for example the doorway scenario: unless either of the green or orange agent alters its course, they will collide. We next give a mathematical definition of SMGs: Beyond saying that SMGs are a subset of general MRN, it is important to highlight the exact characteristics that distinguish the two types of scenarios. We give such a list below.
@@ 2.2 From General MRN to SMGs
2.3 When Does General MRN Become An SMG?
Symmetry arises in MRN when multiple agents occupy equivalent roles, possess similar capabilities, and face identical interaction structures. In symmetric SMGs, such as two robots approaching a narrow doorway, agents have identical incentives and information, resulting in multiple equivalent collision-free solutions [chandra2025deadlock]. Without an explicit symmetry-breaking mechanism, such interactions can converge to undesirable equilibria, including deadlocks or oscillatory behavior.
Geometrically, we quantify symmetry through a liveness function ,
The liveness function is meant to geometrically capture the symmetry of an SMG. The lower the liveness function value, the closer the scenario is to an SMG, and conversely, the higher the liveness function value, the closer the scenario is towards a general MRN. Formally, the liveness function measures the angle between the relative displacement and the relative linear velocity vectors. In the case of perfect symmetry, will be close to as the relative position and velocity vectors are nearly perfectly aligned (parallel) implying the dot product will be nearly . This situation is representative of an SMG where there is perfect contention or conflict over a capacity-one resource such as a doorway or intersection scenario.
The dot product decreases ( increases) as the symmetry decreases. Recall that if there is no symmetry, then there is low chance of a contention or bottleneck. The in the denominator represents the practicality of the fact that there always exists a floating point difference among agents’ velocities, no matter how close they may be.
@@ 2.4 Common SMG Scenarios
In this Section, we briefly list a non-exhaustive set of SMG scenarios that occur in everyday life. These scenarios are compiled from a combination of prior social navigation literature [francis2025principles, francis2023socialhri] and robotics literature [long2018towards, chen2017socially]. Each canonical scenario activates a distinct subset of SMG characteristics. While not exhaustive, they are chosen to cover the most frequently encountered topologies in real life such as doorways, intersections, and blind corners that expose different deadlock mechanisms, visibility challenges, and coordination demands. Additional scenarios, as and when they are discovered and analyzed, will be added to our accompanying software library.
@@ 2.5 Evaluation Metrics
The typical metrics for evaluating general MRN are the following: average , average delay, and path deviation. We chose to include these metrics as they enable benchmarking with general MRN methods. The average metric helps measure movement smoothness, the average delay compares time efficiency between agents, and path deviation assesses navigation accuracy. Together, these metrics ensure that the path planners can handle real-world challenges effectively, navigating efficiently, smoothly, and reliably while avoiding deadlocks. This thorough evaluation is essential for developing robust multi-agent systems in complex environments. Although existing metrics for general MRN remain essential, in an SMG, agents have high agency. We therefore also report the following metrics: flow rate, fairness (game-theoretic perspective), and influence (externalities imposed on others). The following sections discuss each of them in more detail:
Avg (Average Change in Velocity): This metric measures how much an agent’s speed changes on average at each time step during its movement. It shows how often and how much the agent adjusts its speed in response to its surroundings and other agents. Since most real world systems use a discrete-time controller, it is more tractable to use a discrete version of the metric so that the metric clock can be synchronized with the robot clock. The metric is computed as:
@@
where is summed over the entire episode over which the metric is evaluated. A lower Avg means smoother and more consistent movement, with fewer disruptions or sudden changes. A higher Avg means the agent is often changing its speed, which could indicate navigation challenges or the need for constant adjustment in a changing environment. All experiments use the same fixed time-step, so comparisons across methods are well-defined even if the absolute number changes with .
Average Delay: To quantify the temporal inefficiency introduced by multi-agent interactions, we adopt a delay metric inspired by classical traffic flow theory. We measure the additional time incurred by each agent relative to its own nominal travel time in the absence of interactions. For each agent , let denote the time required for the agent to reach its goal in the SMG. We define a baseline travel time as the time required for the same agent to reach the same goal when navigating alone in the same environment under the same planning or control method. The delay experienced by agent is then defined as . Using this definition, we compute the Average Delay across all agents as
This metric captures the average additional time agents experience due to congestion, coordination requirements, or bottleneck contention. A value of zero indicates that agents experience no performance degradation relative to solo navigation. Because nominal travel times may vary across agents and scenarios, we additionally report a normalized version of the metric:
The normalized delay enables fair comparison across heterogeneous agents and environments by measuring relative rather than absolute performance degradation. Larger values indicate greater interaction-induced inefficiency and stronger coupling among agents.
@@
Path Deviation: Path Deviation measures how much an agent’s path differs from a reference or ideal path. It uses the Hausdorff distance, which calculates the greatest distance from a point in one path to the closest point in the other path. Further note that there may exist multiple optimal trajectories. Let denote the set of nominal optimal paths (e.g., straight lines or solutions from a reference planner). We define Path Deviation as
@@
Note that this metric is purposefully kept broad so that it can be adapted to other deviation metrics by the user. In this case, the deviation is measured against the ideal scenario where only one agent moves towards the goal without any interactions or obstacles, defined as the nominal agent scenario. A low Path Deviation means the agent’s actual path is very close to the optimal path, indicating efficient navigation. A high Path Deviation means the agent took significant detours or deviations, possibly due to interactions with other agents, obstacles, avoiding or resolving deadlocks in the way or irregular control generation leading to poor navigation.
Flow Rate: We define the flow rate to quantify how efficiently multiple agents traverse a spatial bottleneck such as a doorway, corridor opening, or narrow passage. Let represent a gap region, defined as a line segment or spatial region (e.g., a doorway cross-section) with physical width (measured in meters). We assume each agent has a convex hull footprint determined by its state . An agent is considered to have passed through the gap when the center of its convex hull transitions from one side of to the other. Let denote the set of agents assigned to traverse the gap . To ensure consistency in evaluation, we define a measurement time interval using a standardized convention. The start time is defined as the earliest time at which any agent in enters a predefined entry region adjacent to . The end time is defined as the time at which the last agent in exits a predefined post-gap region on the opposite side of . Alternative implementations may use a fixed time horizon, but throughout this work we adopt the entry-exit convention described above. The flow rate is then defined as
(2) where denotes the number of agents required to traverse the gap. This metric captures the density-normalized throughput of agents across the bottleneck, and is analogous to classical traffic-flow formulations that measure volume of traffic per unit time normalized by available capacity.
Fairness: We adopt a game-theoretic formulation of fairness that captures how individual agents balance their own objectives with the collective performance of the multi-agent system. Let denote a strategy profile. Let be the trajectory realized by agent when all agents execute the strategy profile . Thus, strategies represent decision-making rules, while trajectories represent the resulting state evolution induced by those rules. Let denote the cumulative reward obtained by agent when all agents execute . We define the global system reward as
(3) where represents the priority or importance weight of agent (e.g., emergency vehicles may be assigned larger values).
While the underlying planning or learning algorithm may seek to optimize either local rewards or the global reward, fairness is evaluated using the realized outcomes induced by the selected strategy profile. Specifically, we measure fairness through the disparity between each agent’s achieved local reward and the reward it would obtain under a globally optimal cooperative strategy:
(4) where denotes a fully cooperative strategy profile and is a small constant used to avoid division by zero.
This metric quantifies the degree to which the executed strategy profile aligns with a socially optimal allocation of rewards. Higher fairness values indicate that agents collectively achieve reward distributions closer to those under the cooperative optimum. Importantly, fairness is evaluated on the realized trajectories induced by the strategy profile rather than on the optimization procedure itself.
Influence Score (IS): The influence score measures the extent to which the presence of an agent alters the behavior of other agents in the system. We define influence using a game-theoretic counterfactual formulation. As before, let denote a strategy profile. For a given agent , we define the counterfactual strategy profile as the strategy profile obtained by removing agent from the environment and re-solving the same SMG using the same planning or learning algorithm for the remaining agents. The resulting control input for agent is denoted .
The influence score of agent is defined as
(5) This metric quantifies how strongly agent perturbs the control behavior of other agents. Higher values of indicate that the presence of agent forces larger control deviations in other agents, suggesting more aggressive or disruptive interaction behavior.
In practice, for deterministic planners or controllers, and are computed using two simulations with identical initial conditions except for removing agent . For stochastic planners or learning-based policies, the control signals may be replaced by expected controls or averaged across multiple simulation rollouts. We emphasize that the influence score is an evaluation metric applied to the realized outcomes induced by a selected strategy profile. It does not modify or redefine the underlying planning or learning objective.